# Revisiting the effects of baryon physics on small-scale redshift space distortions

### v2

@@ Abstract
Redshift space distortions (RSDs) appear as an apparent anisotropy in the clustering of galaxies along the line of sight produced by their peculiar velocities. This motion traces the local gravitational potential and thus observations of the strength of the distortions is a powerful probe of gravity and structure formation. Redshift space distortions are a key probe for many current and planned large volume galaxy surveys such as DESI (Dark Energy Spectroscopic Instrument Collaboration, 2016), Euclid (Laureijs et al., 2011) and the Nancy Grace Roman Space Telescope (WFIRST; Spergel et al., 2013), which aim to measure the linear growth rate of large-scale structure to an accuracy of 2% error. While most measurements involving RSDs focus on the large-scale clustering of galaxies, where the maximum scales used are in the quasi-linear regime of Mpc-1 or Mpc (see, for example, Alam et al., 2017; Satpathy et al., 2017; Gil-Marín et al., 2020; Bautista et al., 2021; de Mattia et al., 2021), analyses using smaller scales in the Mpc range can potentially offer much more constraining power (see, for example, Reid et al., 2014; Wibking et al., 2019; Chapman et al., 2021; Yuan et al., 2022; Zhai et al., 2022; Lange et al., 2023). However, it is important to characterize any possible impact that gas physics and the clustering of baryons may have on the redshift space clustering statistics to remove these effects as a possible source of contamination or systematic bias. Furthermore, many of these studies involving small-scale RSDs have reported measuring lower growth rates, , and clustering amplitudes, than other observations such as the Cosmic Microwave Background (Planck Collaboration, 2020). Indeed, there is a growing volume of literature that has detected small () tensions from the Planck cosmology. In light of these controversies and in planning for future surveys, it becomes timely to examine the systematics that might affect these measurements, particularly as many of the models used to inform parameter estimation from RSD analyses are derived from or tested against gravity only N-body simulations. In particular, we explore how much the redshift space power spectrum is affected by the inclusion of baryonic content and its associated processes.
@@ 2.2.1 Bahamas
The Bahamas (BAryons and HAlos of MAssive Systems) simulations (McCarthy et al., 2017, 2018) are a set of hydrodynamic N-body simulations run using the Smoothed Particle Hydrodynamics (SPH) code GADGET3 SPH (Springel, 2005). Each box has a length of 400 Mpc per side and contains dark matter particles and (initially) SPH particles to represent baryons. The Bahamas simulations can be grouped into roughly two sets: those based on WMAP 9-Yr best fit cosmology (Hinshaw et al., 2013) and those with the Planck 2013 best fit cosmology (Planck Collaboration, 2014). In this work, we only use the simulations in the WMAP9 cosmology, with the following parameters: , , , , and . The feedback models have been carefully calibrated to reproduce the local galaxy stellar mass function (using data from Bernardi et al. 2013; Baldry et al. 2012; Li & White 2009) and the gas mass fractions of galaxy groups and clusters as measured with X-ray observations (Vikhlinin et al., 2006; Maughan et al., 2008; Sun et al., 2009; Pratt et al., 2009; Lin et al., 2012; Lovisari et al., 2015). The Bahamas simulations have been featured in a number of other works that explore the impact of baryons (and massive neutrinos) on large-scale clustering statistics such as the suppression of the real space, matter power spectrum (van Daalen et al., 2020; Acuto et al., 2021; Salcido et al., 2023), halo bias and correlation functions (Pfeifer et al., 2020; Stafford et al., 2020) and the bispectrum (Foreman et al., 2020; Yankelevich et al., 2022).
@@ 3.1 Effect on large scale clustering: Halo mass dependence
Figures 3 and 4 show the changes to the monopole and quadrupole moments in redshift space as a result of the inclusion of baryonic physics. For Figure 3, the mass binning is performed on the total halo mass for each simulation with the halo mass defined as the sum of all the particles belonging to the halo, which in the case of the fiducial AGN simulation includes the gas and stars, as identified by Subfind. Note that because of limited sample numbers, the ratio cannot be interpreted over the entire range shown for M mass haloes. To guide the reader, we have also marked the scale at which the shot noise contribution to the monopole moment of the Bahamas AGN fiducial model becomes dominant. The shot noise has been calculated as , where is the number density of haloes in that mass bin. We can see that there is a substantial difference between the halo-halo power spectra across simulations, even on the very largest scales. This is greatest for haloes in the range 13.0 M 13.5 for which the difference is 7-10% at kMpc in the monopole (3% in the quadrupole) and rises to 13% (15%) on smaller scales. Some of this difference is due to the sample selection, in that the inclusion of baryon has changed the halo abundance so the same haloes are not being compared across the simulations; in fact this very mass range in which we see the greatest difference between the power spectra is also the same range at which the peak difference in the halo mass functions occurs as first noted by van Daalen et al. (2014) (see figure 14 of Pfeifer et al. 2020 for the specific Bahamas mass function). This is because haloes in the 13.0 M 13.5 range in the fiducial AGN simulation are the most susceptible to AGN ejection. Schaller et al. (2015) showed that for the Eagle simulations, the effect of gas physics is to monotonically lower the halo masses except for the very largest clusters at M⊙, meaning that these objects are now rarer and so the halo bias is increased. For low mass haloes of M, the baryons and even the dark matter particles inside the halo are disrupted by feedback processes and the halo radius shrinks, however, as the halo mass increases, stellar feedback becomes less effective and the ratio of halo masses, MAGN/MDMO, approaches the universal dark matter fraction. At even higher masses ( M), the mass ratio approaches unity as the large gravitational potential of the cluster prevents baryon expulsion from AGN feedback. We see a similar effect in Bahamas , except the mass ratio, MAGN/MDMO, does not continue to fall for lower masses, rather there is an upturn at M⊙ and these haloes now have approximately the same halo mass with or without gas physics. This is because the stellar feedback in Eagle is stronger than in Bahamas , but Eagle has weaker AGN feedback. This relationship between MAGN and MDMO is reflected in the monopole moment in Figure LABEL:fig:halo_mass_unmatched, in which both the lowest and highest mass bins exhibit the smallest change in behaviour in the redshift space monopole since they experience the smallest change in mass.
@@
Figure 3 demonstrates that there is a only a percent level difference in the large scale halo bias of both monopole and quadrupole moments if we used matched catalogues to compare the fiducial AGN Bahamas simulation with its dark matter only equivalent. For the lowest mass range, however, a difference of 2.5% remains in the large scale halo bias even for the matched case. We have determined that this is due to missing haloes in the AGN fiducial case causing the matched halo mass function to be systematically lower than the unmatched version by about 2-4% in this mass range, because some of the smallest haloes do not have a dark matter only counterpart. Some of these cases would be due to limitations in the mass resolution in the Bahamas simulations, while others could have undergone a physical process that destroys the halo. Again, while there is no discernible trend with increasing halo mass in the matched quadrupole power spectra (Figure 3, right), there seems to be a steady enhancement of the difference in the quadrupole moments over Mpc-1 up to 15%. This implies that a simple remapping of the halo mass function to its equivalent under gas physics is sufficient to account for large scale changes (Mpc) when baryons are included if we have access to their equivalent haloes that would have evolved under gravity only conditions. We have also checked that this remains true when using a matched sample from the other simulations in the Bahamas suite. Our results are consistent with the trend seen in the real space halo power spectrum (Yankelevich et al., 2022). Furthermore, Yankelevich et al. (2022) show that the (real space) halo bispectrum is more sensitive to these changes in the halo mass, where the halo profile makes an appreciable impact. It would be interesting to see if the same holds true for the redshift space bispectrum, but we leave this for a future exploration.
@@ 3.2.1 Halo mean infall velocity: Halo velocity profile
In Figure 5 we show calculated between central haloes and halo particles of the Bahamas simulations as a function of the distance to the halo center for the total matter content. This is done for four bins in halo mass spanning 0.5 dex (as in Figures LABEL:fig:halo_mass_unmatched and LABEL:fig:halo_mass_matched) for each of the Bahamas simulations listed in Table 1 and a dark matter only counterpart. The top left panel shows the total for all halo particle types and all Bahamas simulations, while the other three panels are ratios with respect to the dark matter only simulation for the fiducial AGN model using a different species each time, namely: all particles (total), gas particles only and star particles only. From the top left panel, we can immediately see that the velocity profiles of the haloes within each mass bin do not seem to differ a great deal across simulations; in fact for clarity, we have omitted the other Bahamas variations when plotting the ratios in the three other panels of Fig 5. Also since no unbinding procedure is performed, some interloper particles and spurious linkage causes measurements of the mean infall velocity to continue to be defined beyond the halo radius. We have checked that the different AGN models only induce a modest 3-4% change relative to the fiducial model in the halo mean radial velocity at most for the gas and stars, rising to 20% at the innermost regions where Mpc. In fact, the mean pairwise velocity for all the Bahamas simulations, seems to agree at every mass scale regardless of the change in halo mass across each simulation; since this is not a matched set of haloes, the mass binning means that a slightly different set of haloes are selected from each simulation. With matched haloes, the differences between the simulations would be reduced even further, but since the differences are quite small to begin with, we have chosen not to show these curves in order to reduce the clutter on this Figure.
@@ 4 Conclusions
We found that contrary to Hellwing et al. (2016), baryonic effects can have a greater than percent level impact on the redshift space matter power spectrum on scales as large as Mpc-1. This is because the Eagle simulation fails to both capture the true extent of gas physics and the transition region between the linear Kaiser model to the non-linear ‘Fingers-of-God’ effect. Furthermore, this effect is scale-dependent and introduces an additional bias between the total matter and galaxy power spectra, as shown in Figure LABEL:fig:halo_mass_unmatched, although the quantitative impact on cosmological parameter estimation has yet to be determined.

### v3

@@ Abstract
Redshift space distortions (RSDs) appear as an apparent anisotropy in the clustering of galaxies along the line of sight produced by their peculiar velocities. This motion traces the local gravitational potential and thus observations of the strength of the distortions is a powerful probe of gravity and structure formation. Redshift space distortions are a key probe for many current and planned large volume galaxy surveys such as DESI (Dark Energy Spectroscopic Instrument Collaboration, 2016), Euclid (Laureijs et al., 2011) and the Nancy Grace Roman Space Telescope (WFIRST; Spergel et al., 2013), which aim to measure the linear growth rate of large-scale structure to an accuracy of 2% error. While most measurements involving RSDs focus on the large-scale clustering of galaxies, where the maximum scales used are in the quasi-linear regime of Mpc-1 or Mpc (see, for example, Alam et al., 2017; Satpathy et al., 2017; Gil-Marín et al., 2020; Bautista et al., 2021; de Mattia et al., 2021), analyses using smaller scales in the Mpc range can potentially offer much more constraining power (see, for example, Reid et al., 2014; Wibking et al., 2019; Chapman et al., 2022; Yuan et al., 2022; Zhai et al., 2023; Lange et al., 2023). However, it is important to characterize any possible impact that gas physics and the clustering of baryons may have on the redshift space clustering statistics to remove these effects as a possible source of contamination or systematic bias. Furthermore, many of these studies involving small-scale RSDs have reported measuring lower growth rates, , and clustering amplitudes, than other observations such as the Cosmic Microwave Background (Planck Collaboration, 2020). Indeed, there is a growing volume of literature that has detected small () tensions from the Planck cosmology. In light of these controversies and in planning for future surveys, it becomes timely to examine the systematics that might affect these measurements, particularly as many of the models used to inform parameter estimation from RSD analyses are derived from or tested against gravity only N-body simulations. In particular, we explore how much the redshift space power spectrum is affected by the inclusion of baryonic content and its associated processes.
@@ 2.2.1 Bahamas
The Bahamas (BAryons and HAlos of MAssive Systems) simulations (McCarthy et al., 2017, 2018) are a set of hydrodynamic N-body simulations run using the Smoothed Particle Hydrodynamics (SPH) code GADGET3 SPH (Springel, 2005). Each box has a length of 400 Mpc per side and contains dark matter particles and (initially) SPH particles to represent baryons. The Bahamas simulations can be grouped into roughly two sets: those based on WMAP 9-Yr best fit cosmology (Hinshaw et al., 2013) and those with the Planck 2013 best fit cosmology (Planck Collaboration, 2014). In this work, we only use the simulations in the WMAP9 cosmology, with the following parameters: , , , , and . The feedback models have been carefully calibrated to reproduce the local galaxy stellar mass function (using data from Bernardi et al. 2013; Baldry et al. 2012; Li & White 2009) and the gas mass fractions of galaxy groups and clusters as measured with X-ray observations (Vikhlinin et al., 2006; Maughan et al., 2008; Sun et al., 2009; Pratt et al., 2009; Lin et al., 2012; Lovisari et al., 2015). The Bahamas simulations have been featured in a number of other works that explore the impact of baryons (and massive neutrinos) on large-scale clustering statistics such as the suppression of the real space, matter power spectrum (van Daalen et al., 2020; Acuto et al., 2021; Salcido et al., 2023), halo bias and correlation functions (Pfeifer et al., 2020; Stafford et al., 2020) and the bispectrum (Foreman et al., 2020; Yankelevich et al., 2023).
@@ 3.1 Effect on large scale clustering: Halo mass dependence
Figures 3 and 4 show the changes to the monopole and quadrupole moments in redshift space as a result of the inclusion of baryonic physics. For Figure 3, the mass binning is performed on the total halo mass for each simulation with the halo mass defined as the sum of all the particles belonging to the halo, which in the case of the fiducial AGN simulation includes the gas and stars, as identified by Subfind. Note that because of limited sample numbers, the ratio cannot be interpreted over the entire range shown for M mass haloes. To guide the reader, we have also marked the scale at which the shot noise contribution to the monopole moment of the Bahamas AGN fiducial model becomes dominant. The shot noise has been calculated as , where is the number density of haloes in that mass bin. We can see that there is a substantial difference between the halo-halo power spectra across simulations, even on the very largest scales. This is greatest for haloes in the range 13.0 M 13.5 for which the difference is 7-10% at kMpc in the monopole (3% in the quadrupole) and rises to 13% (15%) on smaller scales. Some of this difference is due to the sample selection, in that the inclusion of baryon has changed the halo abundance so the same haloes are not being compared across the simulations; in fact this very mass range in which we see the greatest difference between the power spectra is also the same range at which the peak difference in the halo mass functions occurs as first noted by van Daalen et al. (2014) (see figure 14 of Pfeifer et al. 2020 for the specific Bahamas mass function). This is because haloes in the 13.0 M 13.5 range in the fiducial AGN simulation are the most susceptible to AGN ejection. Schaller et al. (2015) showed that for the Eagle simulations, the effect of gas physics is to monotonically lower the halo masses except for the very largest clusters at M⊙, meaning that these objects are now rarer and so the halo bias is increased. For low mass haloes of M, the baryons and even the dark matter particles inside the halo are disrupted by feedback processes and the halo radius shrinks, however, as the halo mass increases, stellar feedback becomes less effective and the ratio of halo masses, MAGN/MDMO, approaches the universal dark matter fraction. At even higher masses ( M), the mass ratio approaches unity as the large gravitational potential of the cluster prevents baryon expulsion from AGN feedback. We see a similar effect in Bahamas , except the mass ratio, MAGN/MDMO, does not continue to fall for lower masses, rather there is an upturn at M⊙ and these haloes now have approximately the same halo mass with or without gas physics. This is because the stellar feedback in Eagle is stronger than in Bahamas , but Eagle has weaker AGN feedback. This relationship between MAGN and MDMO is reflected in the monopole moment in Figure 3, in which both the lowest and highest mass bins exhibit the smallest change in behaviour in the redshift space monopole since they experience the smallest change in mass.
@@
Figure 3 demonstrates that there is a only a percent level difference in the large scale halo bias of both monopole and quadrupole moments if we used matched catalogues to compare the fiducial AGN Bahamas simulation with its dark matter only equivalent. For the lowest mass range, however, a difference of 2.5% remains in the large scale halo bias even for the matched case. We have determined that this is due to missing haloes in the AGN fiducial case causing the matched halo mass function to be systematically lower than the unmatched version by about 2-4% in this mass range, because some of the smallest haloes do not have a dark matter only counterpart. Some of these cases would be due to limitations in the mass resolution in the Bahamas simulations, while others could have undergone a physical process that destroys the halo. Again, while there is no discernible trend with increasing halo mass in the matched quadrupole power spectra (Figure 3, right), there seems to be a steady enhancement of the difference in the quadrupole moments over Mpc-1 up to 15%. This implies that a simple remapping of the halo mass function to its equivalent under gas physics is sufficient to account for large scale changes (Mpc) when baryons are included if we have access to their equivalent haloes that would have evolved under gravity only conditions. We have also checked that this remains true when using a matched sample from the other simulations in the Bahamas suite. Our results are consistent with the trend seen in the real space halo power spectrum (Yankelevich et al., 2023). Furthermore, Yankelevich et al. (2023) show that the (real space) halo bispectrum is more sensitive to these changes in the halo mass, where the halo profile makes an appreciable impact. It would be interesting to see if the same holds true for the redshift space bispectrum, but we leave this for a future exploration.
@@ 3.2.1 Halo mean infall velocity: Halo velocity profile
In Figure 5 we show calculated between central haloes and halo particles of the Bahamas simulations as a function of the distance to the halo center for the total matter content. This is done for four bins in halo mass spanning 0.5 dex (as in Figure 3) for each of the Bahamas simulations listed in Table 1 and a dark matter only counterpart. The top left panel shows the total for all halo particle types and all Bahamas simulations, while the other three panels are ratios with respect to the dark matter only simulation for the fiducial AGN model using a different species each time, namely: all particles (total), gas particles only and star particles only. From the top left panel, we can immediately see that the velocity profiles of the haloes within each mass bin do not seem to differ a great deal across simulations; in fact for clarity, we have omitted the other Bahamas variations when plotting the ratios in the three other panels of Fig 5. Also since no unbinding procedure is performed, some interloper particles and spurious linkage causes measurements of the mean infall velocity to continue to be defined beyond the halo radius. We have checked that the different AGN models only induce a modest 3-4% change relative to the fiducial model in the halo mean radial velocity at most for the gas and stars, rising to 20% at the innermost regions where Mpc. In fact, the mean pairwise velocity for all the Bahamas simulations, seems to agree at every mass scale regardless of the change in halo mass across each simulation; since this is not a matched set of haloes, the mass binning means that a slightly different set of haloes are selected from each simulation. With matched haloes, the differences between the simulations would be reduced even further, but since the differences are quite small to begin with, we have chosen not to show these curves in order to reduce the clutter on this Figure.
@@ 4 Conclusions
We found that contrary to Hellwing et al. (2016), baryonic effects can have a greater than percent level impact on the redshift space matter power spectrum on scales as large as Mpc-1. This is because the Eagle simulation fails to both capture the true extent of gas physics and the transition region between the linear Kaiser model to the non-linear ‘Fingers-of-God’ effect. Furthermore, this effect is scale-dependent and introduces an additional bias between the total matter and galaxy power spectra, as shown in Figure 3, although the quantitative impact on cosmological parameter estimation has yet to be determined.